OpenStudy (anonymous):

Evaluate the following expression: 3(1,-1) - (4,-2). Help me please

OpenStudy (ipwnbunnies):

Are these vectors?

OpenStudy (anonymous):

Yeah haha ipwnbunnies sick name manXD

OpenStudy (ipwnbunnies):

Apply the scalar multiple to each component of the vector: \[a<x,y> = <ax,ay>\]

OpenStudy (ipwnbunnies):

When adding or subtracting vectors, add or subtract the similar components to get the resultant vector. As in the x-components add and subtract from other x-components, etc.

OpenStudy (anonymous):

Could you possibly show me how to do that sorry if I horrible at this I just want to learn

OpenStudy (ipwnbunnies):

Let's say we have this problem: \[3<4,-3>\]

OpenStudy (anonymous):

Alright

OpenStudy (ipwnbunnies):

We multiply both components of the vector by the scalar number, the number outside. So, simplifying this problem will give us: \[<12,-9>\]

OpenStudy (anonymous):

Alright cool man but that isn't one of the options on here?

OpenStudy (ipwnbunnies):

That was an example...

OpenStudy (anonymous):

Oh okay my b

OpenStudy (ipwnbunnies):

Do the same for the one vector in your problem, then add the vectors.

OpenStudy (anonymous):

Okay so it'll be 3-3+12+6?

OpenStudy (ipwnbunnies):

No. So we have these two vectors now: \[<3,-3> - <4,-2>\]

OpenStudy (anonymous):

Okay

OpenStudy (ipwnbunnies):

You subtract the similar components. As in the x-components and subtracted, the y-components are subtracted, etc.

OpenStudy (ipwnbunnies):

The result is a vector, your final answer.

OpenStudy (anonymous):

So<-1,5>?

OpenStudy (ipwnbunnies):

Try again. You did some bad math, lol. \[<3-4,-3-(-2)>\]

OpenStudy (anonymous):

<-1-6> right?

OpenStudy (ipwnbunnies):

...No. Do you know how to subtract negative numbers? -2 - (-4) = -2 = 4 = 2

OpenStudy (anonymous):

-2-(-4)=8? Cause you distribute the negative sorry man I'm confused

OpenStudy (ipwnbunnies):

Yeah, you distribute the negative sign, but it is still an addition or subtraction problem. -2 - (-4) Multiply -4 by -1: -2 + 4 = 2

OpenStudy (anonymous):

Oh okay my bad but 2 isn't an option either :/

OpenStudy (ipwnbunnies):

It's not 2 either... \[<3-4, -3 - (-2)>\] \[<-1,-1>\]

OpenStudy (ipwnbunnies):

-3 - (-2) = -3 + 2 = -1

OpenStudy (anonymous):

Thanks man sorry about all that I'm math challenged